OpenStudy (anonymous):

\[\frac{ 4x-2 }{ 3 }+\frac{ 2x-8 }{ 2 }=5\]

OpenStudy (anonymous):

Really need help.

OpenStudy (mertsj):

1. Multiply both sides by 6 2. Collect the like terms 3. Add 28 to both sides. 4. Divide both sides by 14

OpenStudy (anonymous):

Okay, all the sides by 6?

OpenStudy (mertsj):

Every term by 6

OpenStudy (anonymous):

Is there any way you could show me, i'm sorry i'm really confused

OpenStudy (mertsj):

\[6\times \frac{4x-2}{3}+6 \times\frac{2x-8}{2}= 5 \times 6\]